OpenStudy (4meisu):

simplify 9y - [4y-(2y-5)]

OpenStudy (lalaly):

\[9y-4y-2y+5 = 3y+5\]

OpenStudy (4meisu):

Where did the 3y come from?

OpenStudy (lalaly):

9y-4y-2y=3y

OpenStudy (4meisu):

Wouldn't you have to multiply 4y and 2y though? Since 2y-5 is in a bracket

OpenStudy (anonymous):

9y - [4y-(2y-5)] You break [4y-(2y-5)] apart by distributing the - that is in front of it -[4y-(2y-5)] same thing with -(2z-5) once you break that apart you have 9y-4y-2y+5 = 3y+5

OpenStudy (4meisu):

thanks guys!